OpenStudy (anonymous):

let f(x)=-4x+7 and g(x)=10x-6 find f(g(x))

OpenStudy (ivettef365):

you are missing something, is it f(g(x)) ?

OpenStudy (anonymous):

oh yea!! im sorry your right

OpenStudy (debbieg):

So plug g(x) into f(x) everywhere there is an x, then simplify. What do you get?

OpenStudy (anonymous):

oops was the answer wrong?

OpenStudy (anonymous):

well the answer i got when i finished working out the problem is 4

OpenStudy (anonymous):

@satellite73 no this is diffrent :)

OpenStudy (anonymous):

oh ok but it looks identical

OpenStudy (anonymous):

wow i put up the wrong one againn!!!

OpenStudy (debbieg):

Your answer should be a function of x.

OpenStudy (anonymous):

its really let f(x)=2x+2 solve f^1(x) when x=4

OpenStudy (anonymous):

and i got 4 for my answer

OpenStudy (debbieg):

So you need \[\large f ^{-1}(4)\]?

OpenStudy (anonymous):

no \[f^1(x)\]

OpenStudy (anonymous):

i was saying my answer for the whole equation is 4 but i wasnt sure if im right

OpenStudy (debbieg):

\[\large f\prime(x)\]?

OpenStudy (anonymous):

i just showed you what it was above

OpenStudy (debbieg):

well, I don't know what you mean by \[\large f ^{1}(x)\]other than just \[\large f(x)\]. That notation does not make sense. if it were \[\large f\prime(x)\]or \[\large f ^{-1}(x)\]that would make sense. But as far as I can tell, \[\large f ^{1}(x)=f(x)\]so I guess at x=4 that would be 2(4)+2=10.